The sum of two numbers is 240.The larger number is 6 less than twice the smaller. find the numbers
Answer by piglet162431(90)

Let n be the Smaller number
n + 2n-6=240
3n-6=240
3n=246
n=82
if the smaller number is 82, then 82(2)-6=164-6=158
Check.
158+82=240